10. A pulley system operated with an electric motor consumes 5.0 kJ of energy as it does 3.2 kJ of work in raising an object. The efficiency of this technology is:

(a) 64% (b) 84%

(c) 119% (d) 156%



1
Expert's answer
2022-04-17T17:07:33-0400

By definition, the efficiency is given as follows:


"\eta = \dfrac{W}{E}\times 100\%"

where "W = 3.2 kJ" is the produced work, and "E = 5.0 kJ" is the consumed energy. Thus, obtain:


"\eta = \dfrac{3.2 kJ}{5.0 kJ}\times 100\% =64\%"

Answer. (a) 64%.